if 20 men do a piece of work in 8 days .how many men will do a work in 10 days ?
1
Expert's answer
2020-02-10T09:31:29-0500

The speed of 1 worker:


"\frac{1}{20\cdot 8}=\frac{1}{160}\frac{1}{day}"

The number of workers will be:


"\frac{1}{\frac{1}{160}10}=16\"